OpenStudy (anonymous):

What is the slope of a line parallel to the line whose equation is 2x – 3y = 6?

OpenStudy (anonymous):

Write the equation in terms of y first.

OpenStudy (anonymous):

how do I do that..?

OpenStudy (anonymous):

2x - 3y = 6 -2x -2x -3y = -2x+6 divide all sides by -3 y = (2/3)x-2

OpenStudy (anonymous):

The number next to the x (2/3) is the slope

OpenStudy (anonymous):

none of this makes sense to me..im so stupid

OpenStudy (anonymous):

A line which is parallel to another line has the same slope.

OpenStudy (anonymous):

To solve for y, you get y on one side, and everything else on the other side

OpenStudy (anonymous):

If the slope is the same, let x =(2/3), and the following number something else other than -2

OpenStudy (anonymous):

y = (2/3)x+5 is parallel to y=(2/3)x-2